Solve for the t-computed value of the following. Write your answer to the nearest

hundredths. Show the complete solution.


  1. "\chi" =15 "\mu" = 12 s = 4 n = 12
  2. "\chi" = 7.8 "\mu" = 10 s = 5.1 n = 10
  3. "\chi" = 5 "\mu" = 5.3 s = 0.14 n = 8
  4. "\chi" = 7.2 "\mu" = 9.1 s = 1.8 n = 15
  5. "\chi" = 5 "\mu" = 7 s = 3 n = 10
1
Expert's answer
2022-05-04T13:05:07-0400

1.


"t=\dfrac{x-\mu}{s/\sqrt{n}}=\dfrac{15-12}{4/\sqrt{12}}\approx2.60"

2.


"t=\dfrac{x-\mu}{s/\sqrt{n}}=\dfrac{7.8-10}{5.1/\sqrt{10}}\approx-1.36"

3.


"t=\dfrac{x-\mu}{s/\sqrt{n}}=\dfrac{5-5.3}{0.14/\sqrt{8}}\approx-6.06"

4.


"t=\dfrac{x-\mu}{s/\sqrt{n}}=\dfrac{7.2-9.1}{1.8/\sqrt{15}}\approx-4.09"

5.


"t=\dfrac{x-\mu}{s/\sqrt{n}}=\dfrac{5-7}{3/\sqrt{10}}\approx-2.11"